Refer to the following distribution of commissions:

Monthly CommissionsClass Frequencies

$600 up to $8003

$800 up to $1,0007

$1,000 up to $1,20011

$1,200 up to $1,40022

$1,400 up to $1,60040

$1,600 up to $1,80024

$1,800 up to $2,0009

$2,000 up to $2,2004

What is the relative frequency of salespeople who earn $1,600 or more?
Answer #1

Monthly Commission

Class

Frequency Relative frequency
600 - 800 3 3/120
800 - 1000 7 7/120
1000 - 1200 11 11/120
1200 - 1400 22 22/120
1400 - 1600 40 40/120
1600 - 1800 24 24/120
1800 - 2000 9 9/120
2000 - 2200 4 4/120
Total 120

P(X>= 1600) = (24/120) + (9/120) + (4/120)

= (24+9+4)/120

= 37/120

= 0.3083

relative frequency of salespeople who earn $1,600 or more is (37/120).